# Investigating Bias in LLM-Based Bias Detection: Disparities between LLMs and Human Perception

### v1

@@ Abstract
The pervasive spread of misinformation and disinformation in social media underscores the critical importance of detecting media bias. While robust Large Language Models (LLMs) have emerged as foundational tools for bias prediction, concerns about inherent biases within these models persist. In this work, we investigate the presence and nature of bias within LLMs and its consequential impact on media bias detection. Departing from conventional approaches that focus solely on bias detection in media content, we delve into biases within the LLM systems themselves. Through meticulous examination, we probe whether LLMs exhibit biases, particularly in political bias prediction and text continuation tasks. Additionally, we explore bias across diverse topics, aiming to uncover nuanced variations in bias expression within the LLM framework. Importantly, we propose debiasing strategies, including prompt engineering and model fine-tuning. Extensive analysis of bias tendencies across different LLMs sheds light on the broader landscape of bias propagation in language models. This study advances our understanding of LLM bias, offering critical insights into its implications for bias detection tasks and paving the way for more robust and equitable AI systems.
@@
[…]
@@ 1 Introduction
Detecting media bias Yu et al. (2008); Iyyer et al. (2014); Liu et al. (2022) was crucial due to the pervasive spread of misinformation and disinformation on social media platforms, profoundly shaping public perception and decision-making processes. Recently, researchers have increasingly turned to robust LLMs as foundational tools for media bias prediction Lin et al. (2024a, b); Liu et al. (2024). Compared to non-pretrained neural models or less powerful language models, LLMs offer enhanced capabilities, yet with an increased risk of bias introduction, given their superior performance and widespread use in media analysis and bias detection. Consequently, there is a growing need to examine bias within the bias detection process itself Fang et al. (2023); Urman and Makhortykh (2023); Esiobu et al. (2023).
@@ 3.1 LLM-based Bias Prediction
Dataset
Bias Label # of Instances Avg Length
Source
@@
We adopt vanilla ChatGPT model to conduct political leaning prediction on two popular datasets (i.e., FlipBias Chen et al. (2018) and ABP Baly et al. (2020)). The statistic of these two datasets can be found in Table 1. We can see that there are 1022 triples (i.e., each triple is with left-, center-, right-leaning article on same event) in FlipBias and more than 30k instances in ABP dataset. For each instance, we prompt gpt-3.5-turbo-0613 with the following instruction to get the bias prediction results of vanilla ChatGPT:
@@ 3.2 LLM-based Article Continuation
We instruct gpt-3.5-turbo-0613 with a simple article continuation prompt, as follows, to generate a continuation based on the provided prefix.
@@
By following Yano et al. (2010), we first construct two vocabularies for left-leaning and right-leaning articles separately. Each vocabulary is constructed by doing statistic of the word frequency for articles with ground-truth left and right labels and removing stop words (details can be found in Appendix A), which can represent the characteristic of the respective political leaning. The presence of a higher number of words from a specific vocabulary within an article indicates the alignment of the article with the corresponding political leaning. For instance, an article featuring more tokens from the left-leaning vocabulary indicates its left-leaning orientation.
@@
In Fig. 3, we present the outcomes of article continuation experiments with varying prefix lengths (e.g., 20, 40 tokens) employing both embedding-based and vocabulary-based matching. It’s important to note that only the relative percentages of left and right are presented, disregarding the center situation. From Fig. 3, we can see that across prefix lengths ranging from 20 to 80, both label matching methods consistently show a higher percentage of left predictions, suggesting a left-leaning trend in continued articles. However, as the prefix length increases to 320, both methods begin to predict continued articles as more right-leaning. This change may be attributed to the fact that the average length of right-leaning articles is obviously shorter than left-leaning articles. Therefore, when given a prefix with 320 tokens, the leaning of the prefix becomes clear, allowing the suffix generated by the LLM to be successfully predicted as right-leaning. Taking into account these factors, we can conclude that the examined LLM tends to lean more towards the left, as its continued articles are predominantly identified as left-leaning rather than right-leaning.
@@ 4 RQ2: Do LLMs demonstrate consistent bias across all topics?
As elaborated in Section 3.1, our tested LLM exhibits a left-leaning bias compared to viewpoints derived from the ground-truth labels assigned by human evaluators. In this section, we delve into whether the LLM consistently showcases a leaning across all discussed topics. While the ABP dataset includes topic information, the FlipBias dataset lacks such information inherently. To address this, we construct latent topics following the methodology proposed by Lin et al. (2024b). The detailed process of latent topic construction is provided in Appendix B. As the constructed latent topics of FlipBias dataset are not predefined, we attempt to demonstrate their relevance and coherence to predefined topics in ABP dataset. This is achieved by presenting statistics on the (latent) topics of both datasets in Table 2, and by plotting joint distributions of Left-Center (i.e., where the ground-truth label is left and the predicted label is center) and Right-Center, as well as Center-Left and Center-Right, accompanied by kernel density estimates in Fig. 4. It is evident that the joint plots of the FlipBias and ABP datasets exhibit similar patterns. The main difference arises in the distributions based on predefined topics (i.e., Fig. 4(c) and Fig. 4(d)), which appear more focused compared to the distributions based on latent topics (i.e., Fig. 4(a) and Fig. 4(b)), which demonstrate greater dispersion.
@@
Interpretation of Top and Bottom 5 latent topics (ranked by BTI-1 values)
BTI-1 BTI-2 Frequency
Comprehensive Use of Quotes and Citations in Journalism
0.44 0.00 81
Diverse Perspectives on President Obama’s Policies and Actions
0.31 0.00 99
Analysis of Recent Terrorist Attacks and Security Measures in Various Cities
0.31 0.00 103
Critique of DACA Amnesty Program and Advocacy for Stricter Immigration Policies
0.29 0.00 89
Diverse Rhetorical Strategies in Political Discourse
0.29 0.00 97
Trump’s Clashes with Federal Law Enforcement and Media
-0.17 0.11 80
Analysis of Media Coverage Surrounding Trump’s Ratings, Criticisms, and Mental Fitness
-0.20 -0.03 98
Trump Administration’s Response to Russia Sanctions and Political Fallout
-0.27 0.04 71
Satirical Commentary and Critique on Political Events and Figures
-0.30 0.00 58
Media Coverage of Trump Administration
-0.58 0.02 70
Table 3: Interpretation of Top and Bottom 5 Latent Topics on FlipBias.
@@
From Table 3, we observe that the trend of BTI-2 values is more centered around $0.0$ when the range of BTI-2 extends to ±$\pm 0.5$, which is consistent with the observation of Fig. 5. The language model’s left-leaning viewpoints on topics (upper part of Table 3) like journalism’s use of citations, Obama’s policies, and immigration critique reflect values of transparency, inclusivity, and social justice. This aligns with the narrative often seen in left-leaning media, emphasizing fact-checking, diverse perspectives, and human rights advocacy. These viewpoints may shaped by the model’s training data and structural biases. The prevalence of Trump-related topics among the bottom 5 latent topics (lower part of Table 3) with negative BTI-1 suggests a potential right-leaning bias in the language model’s treatment of Trump administration subjects. Given FlipBias’s data collection primarily from 2013 to 2018, a period marked by heightened political polarization, this alignment hints at a correlation between temporal context and exhibited biases.
@@ 5.2 Fine-Tuning to Debias
By observing the results of Fig. 2, we infer that the LLM demonstrates better performance in clarifying right-label articles from center-label articles compared to clarifying left-label articles from center-label ones. This observation suggests a potential deficiency in the LLM’s ability to accurately recognize left-leaning evaluation criteria. To address this, we adjust the proportion of left-leaning articles in the fine-tuning instances to investigate how varying proportions impact the debiasing process. Specifically, we fine-tune gpt-3.5-turbo using 300 labeled instances with three different proportions: all left-label articles (L-FT), a mixture of left-label and center-label articles (LC-FT), and an equal distribution of left-label, center-label, and right-label articles (LCR-FT).
@@ 5.3 Assessment of Debiasing Strategies
We evaluate the debiasing methods introduced in Section 5.1 and Section 5.2 in this subsection.
@@
The debiasing results on FlipBias are reported in Table 4 and the evaluration metrics are following Lin et al. (2024b). We observe that while finetuning methods generally exhibit better bias prediction performance gains (e.g., better BiF1 and MaF1), they also introduce more bias to the finetuned LLMs, as reflected by larger BTI-1 or BTI-2 values after finetuning. On the other hand, prompt-based debiasing methods show impressive effects, especially DS Ganguli et al. (2023), which is extremely easy yet effective.
@@
We further display the bias tendency index (BTI) distribution on FlipBias after applying some representative debiasing methods in Fig. 7, while distributions of additional debiasing methods and results from the ABP dataset can be found in Section C.2. From Fig. 7, we observe that prompt engineering-based debiasing shows better results, as reflected in the BTI values for topics being centered around $0.0$, which is consistent with the general performance comparison results we introduced in the last paragraph. Additionally, the overall shift in the BTI distribution after LCR-FT debiasing, as shown in Fig. 7(d), indicates that finetuning LLMs may result in better performance (refer to bias prediction results reported in Table 4), but it may inadvertently introduce more severe bias.
@@ 6 RQ4: Do various LLMs exhibit similar bias tendencies?
Figure 8: BTI distribution of Topics on FlipBias dataset for various LLMs.
@@ Appendix A Left-Right Vocabulary Corpus Construction
We use the dataset ABP to construct the Left-Right vocabulary corpus. In the preprocessing step, we use the Python package nltk. First, all articles of ABP are tokenized to collect vocabulary, and then they are converted to lowercase and filtered by the stopwords corpus. The tokens are labeled according to the labels of the articles they belong to. When constructing the Left-Right Vocabulary Corpus, for tokens with both Left and Right labels, we only consider the tokens labeled with much higher frequencies. For example, we calculate the Left ratio by dividing the frequency of this token appearing in Left articles by the sum of all tokens in Left articles. Similarly, we calculate the Right ratio. Only if the Left ratio is more than twice the Right ratio, it will be classified into the Left vocabulary list.
Finally, we select the top 2000 most frequent tokens from the Left vocabulary list, and then select the top tokens from the Right vocabulary list with the same total frequency sum as in the Left, which is 1295 tokens, to build the Left-Right Vocabulary Corpus. We also verify this using the vocabulary of Yano et al. (2010). The constructed vocabularies will be made publicly available for future research.

### v2

@@ Abstract
The pervasive spread of misinformation and disinformation in social media underscores the critical importance of detecting media bias. While robust Large Language Models (LLMs) have emerged as foundational tools for bias prediction, concerns about inherent biases within these models persist. In this work, we investigate the presence and nature of bias within LLMs and its consequential impact on media bias detection. Departing from conventional approaches that focus solely on bias detection in media content, we delve into biases within the LLM systems themselves. Through meticulous examination, we probe whether LLMs exhibit biases, particularly in political bias prediction and text continuation tasks. Additionally, we explore bias across diverse topics, aiming to uncover nuanced variations in bias expression within the LLM framework. Importantly, we propose debiasing strategies, including prompt engineering and model fine-tuning. Extensive analysis of bias tendencies across different LLMs sheds light on the broader landscape of bias propagation in language models. This study advances our understanding of LLM bias, offering critical insights into its implications for bias detection tasks and paving the way for more robust and equitable AI systems [^1] The code is available at https://github.com/lylin0/lin2024investigating.
@@
Luyang Lin1,2, Lingzhi Wang3thanks:   Lingzhi Wang is the corresponding author., Jinsong Guo4, Kam-Fai Wong1,2 1The Chinese University of Hong Kong, China 2MoE Key Laboratory of High Confidence Software Technologies, China 3Harbin Institute of Technology, Shenzhen, China  4University College London, UK 1,2{lylin, kfwong}@se.cuhk.edu.hk 3wanglingzhi@hit.edu.cn  4jinsong.guo@ucl.ac.uk
@@ 1 Introduction
Detecting media bias Yu et al. (2008); Iyyer et al. (2014); Liu et al. (2022) was crucial due to the pervasive spread of misinformation and disinformation on social media platforms, profoundly shaping public perception and decision-making processes. Recently, researchers have increasingly turned to robust LLMs as foundational tools for media bias prediction Lin et al. ; Lin et al. (2024); Liu et al. (2024). Compared to non-pretrained neural models or less powerful language models, LLMs offer enhanced capabilities, yet with an increased risk of bias introduction, given their superior performance and widespread use in media analysis and bias detection. Consequently, there is a growing need to examine bias within the bias detection process itself Fang et al. (2023); Urman and Makhortykh (2023); Esiobu et al. (2023).
@@ 3.1 LLM-based Bias Prediction
Dataset Bias Label # of Instances Avg Length Source
@@
We adopt vanilla ChatGPT model to conduct political leaning prediction on two popular datasets (i.e., FlipBias Chen et al. (2018) and ABP Baly et al. (2020)). The statistic of these two datasets can be found in Table 1. We can see that there are 1022 triples (i.e., each triple is with left-, center-, right-leaning article on the same event) in FlipBias and more than 30k instances in ABP dataset. For each instance, we prompt gpt-3.5-turbo-0613 with the following instruction to get the bias prediction results of vanilla ChatGPT:
@@ 3.2 LLM-based Article Continuation
We prompt gpt-3.5-turbo-0613 with a continuation prompt to generate text based on the given prefix.
@@
By following Yano et al. (2010), we first construct two vocabularies for left-and right-leaning articles separately. Each vocabulary is constructed by doing statistic of the word frequency for articles with ground-truth left and right labels and removing stop words (details are shown in Appendix A), which can represent the characteristic of the respective political leaning. The presence of a higher number of words from a specific vocabulary within an article indicates the alignment of the article with the corresponding political leaning. For instance, an article featuring more tokens from the left-leaning vocabulary indicates its left-leaning orientation.
@@
In Fig. 4, we present the outcomes of article continuation experiments with varying prefix lengths (e.g., 20, 40 tokens) employing both embedding-based and vocabulary-based matching. It’s important to note that only the relative percentages of left and right are presented, disregarding the center situation. From Fig. 4, we can see that across prefix lengths ranging from 20 to 80, both label matching methods consistently show a higher percentage of left predictions, suggesting a left-leaning trend in continued articles. However, as the prefix length increases to 320, both methods begin to predict continued articles as more right-leaning. This change may be attributed to the fact that the average length of right-leaning articles is shorter than left-leaning articles (refer to Fig. 2). Additionally, an analysis of left-leaning articles reveals that sentences with left-leaning bias typically appear in the latter part of the article and rarely appear in the prefix. Therefore, when given a prefix with 320 tokens, the political leaning of the prefix becomes clearer, representing a substantial portion—approximately 40%—of the average length of Right articles (794 tokens) and 28% of Left articles ($1$ tokens). This clearer representation of political leaning in the prefix makes it more likely for the LLM to generate a right-leaning suffix. Consequently, LLMs may find it easier to predict right-leaning continued suffixes.
Discussion about Results of Classifier
We also conduct the experiments based on the classifier gpt-3.5-turbo-0613. We observe that when we use the vanilla ChatGPT model, the model predicts about $70\%$ of the generated articles as Center, compared to around $15\%$ of the articles matched as Center in our other methods. The model prefers to regard the text generated by itself as objective. It needs to be further considered whether we can use a classifier that may contain bias to measure the bias. When we give the model three articles of the same event as references (the three articles are the same as we use in embedding-based similarity matching), the trends are similar to the results of embedding-based similarity. More details and results are in Section E.3.
@@ 4 RQ2: Do LLMs demonstrate consistent bias across all topics?
As elaborated in Section 3.1, our tested LLM exhibits a left-leaning bias compared to viewpoints derived from the ground-truth labels assigned by human evaluators. In this section, we delve into whether the LLM consistently showcases a leaning across all discussed topics. While the ABP dataset includes topic information, the FlipBias dataset lacks such information inherently. To address this, we construct latent topics following the methodology proposed by Lin et al. (2024). The detailed process of latent topic construction is provided in Appendix B. As the constructed latent topics of FlipBias dataset are not predefined, we attempt to demonstrate their relevance and coherence to predefined topics in ABP dataset. This is achieved by presenting statistics on the (latent) topics of both datasets in Table 2, and by plotting joint distributions of Left-Center (i.e., where the ground-truth label is left and the predicted label is center) and Right-Center, as well as Center-Left and Center-Right, accompanied by kernel density estimates in Fig. 5. To interpret the figure, points farther from the origin indicate more incorrect judgments, while deviation from the $x=y$ line reflects an imbalance between cases classified as Right and Left. It is evident that the joint plots of the FlipBias and ABP datasets exhibit similar patterns. The main difference arises in the distributions based on predefined topics (i.e., Fig. 5(c) and Fig. 5(d)), which appear more focused compared to the distributions based on latent topics (i.e., Fig. 5(a) and Fig. 5(b)), which demonstrate greater dispersion.
@@
Interpretation of Top and Bottom 5 latent topics (ranked by BTI-1 values) BTI-1 BTI-2 Frequency
Comprehensive Use of Quotes and Citations in Journalism 0.44 0.00 81
Diverse Perspectives on President Obama’s Policies and Actions 0.31 0.00 99
Analysis of Recent Terrorist Attacks and Security Measures in Various Cities 0.31 0.00 103
Critique of DACA Amnesty Program and Advocacy for Stricter Immigration Policies 0.29 0.00 89
Diverse Rhetorical Strategies in Political Discourse 0.29 0.00 97
Trump’s Clashes with Federal Law Enforcement and Media -0.17 0.11 80
Analysis of Media Coverage Surrounding Trump’s Ratings, Criticisms, and Mental Fitness -0.20 -0.03 98
Trump Administration’s Response to Russia Sanctions and Political Fallout -0.27 0.04 71
Satirical Commentary and Critique on Political Events and Figures -0.30 0.00 58
Media Coverage of Trump Administration -0.58 0.02 70
Table 3: Interpretation of Top and Bottom 5 Latent Topics on FlipBias. When interpreting BTI-1 and BTI-2 values, a BTI value closer to zero indicates less bias in the model. Positive values suggest a left-leaning bias, while negative values indicate a right-leaning bias.
@@
From Table 3, we observe that the trend of BTI-2 values is more centered around $0.0$ when the range of BTI-2 extends to ±$\pm 0.5$, which is consistent with the observation of Fig. 6. The LLM’s left-leaning viewpoints on topics (upper part of Table 3) like journalism’s use of citations, Obama’s policies, and immigration critique reflect values of transparency, inclusivity, and social justice. This aligns with the narrative often seen in left-leaning media, emphasizing fact-checking, diverse perspectives, and human rights advocacy. These viewpoints may shaped by the model’s training data and structural biases. The prevalence of Trump-related topics among the bottom 5 latent topics (lower part of Table 3) with negative BTI-1 suggests a potential right-leaning bias in the language model’s treatment of Trump administration subjects. Given FlipBias’s data collection primarily from 2013 to 2018, a period marked by heightened political polarization, this alignment hints at a correlation between temporal context and exhibited biases.
@@ 5.2 Fine-Tuning to Debias
By observing the results of Fig. 3, we infer that the LLM demonstrates better performance in clarifying right-label articles from center-label articles compared to clarifying left-label articles from center-label ones. This observation suggests a potential deficiency in the LLM’s ability to accurately recognize left-leaning evaluation criteria. To address this, we adjust the proportion of left-leaning articles in the fine-tuning instances to investigate how varying proportions impact the debiasing process. Specifically, we fine-tune gpt-3.5-turbo using 300 labeled instances (sampled from the regular training sets of datasets) with three different proportions: all left-label articles (L-FT), a mixture of left-label and center-label articles (LC-FT), and an equal distribution of left-label, center-label, and right-label articles (LCR-FT). We also conduct the LCR-FT setting on the baseline BERT (pre-trained model bert-base-cased) for a better comparison.
@@ 5.3 Assessment of Debiasing Strategies
We evaluate the debiasing methods introduced in Section 5.1 and Section 5.2 in this subsection. Apart from BTI, the other metrics follow Lin et al. (2024).
@@
The debiasing results on FlipBias are reported in Table 4. We observe that while finetuning methods generally exhibit better bias prediction performance gains (e.g., better BiF1 and MaF1), they also introduce more bias to the finetuned LLMs, as reflected by larger BTI-1 or BTI-2 values after finetuning. On the other hand, prompt-based debiasing methods show impressive effects, especially DS Ganguli et al. (2023), which is extremely easy yet effective.
@@
We further display the bias tendency index (BTI) distribution on FlipBias after applying some representative debiasing methods in Fig. 8, while distributions of additional debiasing methods and results from the ABP dataset can be found in Appendix D. From Fig. 8, we observe that prompt engineering-based debiasing shows better results, as reflected in the BTI values for topics being centered around $0.0$, which is consistent with the general performance comparison results we introduced in the last paragraph. Additionally, the overall shift in the BTI distribution after LCR-FT debiasing, as shown in Fig. 8(d), indicates that finetuning LLMs may result in better performance (refer to bias prediction results reported in Table 4), but it may inadvertently introduce more severe bias.
In conclusion, our debiasing experiments reveal that the strategy Bias Label Explanation and Debiasing Statements make the model more cautious in making biased judgments, albeit at the cost of accuracy. In contrast, few-shot and fine-tuning methods yield better performance but are less effective at mitigating bias
@@ 6 RQ4: Do various LLMs exhibit similar bias tendencies?
Figure 9: BTI of Topics on FlipBias for various LLMs.
@@ 7 Conclusion
Acknowledgments
This research work is partially supported by CUHK direct grant No. 4055209 and CUHK Knowledge Transfer Project Fund No. KPF23GWP20. We are also grateful to the anonymous reviewers for their comments.
@@ Appendix A Left-Right Vocabulary Corpus Construction
We construct the Left-Right vocabulary corpus using the ABP dataset. Initially, all articles in ABP are tokenized using the NLTK Python package. Tokens are converted to lowercase and filtered using a stopwords corpus. Each token is then labeled based on the articles they appear in.
To create the Left-Right Vocabulary Corpus, we prioritize tokens labeled with significantly higher frequencies in either Left or Right articles. Specifically, we calculate the Left ratio by dividing a token’s frequency in Left articles by the total tokens in Left articles, and similarly for the Right ratio. Tokens are included in the Left vocabulary list only if the Left ratio is more than twice the Right ratio.
From the Left vocabulary list, we select the top 2000 most frequent tokens. We then select 1295 tokens from the Right vocabulary list to match the total frequency sum of the Left tokens. This corpus is validated against the vocabulary of Yano et al. (2010). The constructed vocabularies will be publicly available for future research.
Indicators Interpreted Topic
"Provides figures and quotes from individuals involved in the issue", "The article cites statements from various individuals involved in the case, including lawyers, politicians, and advocacy groups.", "The text quotes various experts and government officials to support its claims.", "cites tweets and quotes from Trump, experts, and state officials to support the claims made", "Quotes from various food experts and diplomats.", "The article cites multiple sources, including government documents and quotes from officials." … … Comprehensive Use of Quotes and Citations in Journalism
"Describes President Obama’s decision as "benighted" and "cowardly" while praising President Trump’s decision", ""swipes at Joe Biden," "knocks primary rival Bernie Sanders," "gripes about former President Barack Obama"", "frames the issue as a result of understaffing and mismanagement, blames the Obama administration, and highlights the need to protect the president", "Celebratory tone towards Obama, sarcastic and mocking tone towards Democrats", "Portrays Democrats as wanting a grander celebration, mocks Obama and the holiday", "Frames the decision as a potential unwinding of an Obama executive action, includes criticism from Democrats and environmental groups", "Describes the tough-on-crime approach as a reversal of Obama’s "Smart on Crime" policy, implying a negative change" … … Diverse Perspectives on President Obama’s Policies and Actions
"Mentions celebrations and security measures in various cities", "Mentions specific incidents of terrorism and security measures in different cities", "Presents the incident as a terrorist attack and highlights the victims’ nationalities", "Mentions previous vehicle attacks and quotes from witnesses", "Provides examples of other major music event attacks", "Mentions the arsenal of weapons and ammunition recovered, suggesting the possibility of an accomplice", "Mentions the London subway station fire as a terrorist incident caused by an improvised explosive device", "The article provides examples of previous attacks and the use of improvised explosive devices." … … Analysis of Recent Terrorist Attacks and Security Measures in Various Cities
Table 6: Clustered Indicators and Interpreted Topics.
Interpretation of Top and Bottom 5 latent topics (ranked by BTI-2 values) BTI-1 BTI-2 Frequency
Trump’s Clashes with Federal Law Enforcement and Media -0.17 0.11 80
Examining Controversial Tactics: Dissecting Allegations and Defenses in Recent Political Affairs 0.18 0.10 64
Analysis of Congressional Dynamics: Trump’s Strategy, Witness Battles, and Financial Focus -0.05 0.09 72
Bipartisan Cooperation in Senate: Struggles and Progress -0.05 0.08 68
Unveiling the Constitutional Crisis: Examining Government Overreach and the Erosion of Rights 0.25 0.07 72
Understanding Textual Analysis: The Importance of Examples and Analogies 0.18 -0.05 59
Statewide Controversies: Voter Rights, Criminal Justice, and Transition Integrity 0.18 -0.06 60
Analyzing Political Discourse: Insights from Trump Administration and Beyond -0.02 -0.08 76
Examining Biased Reporting in Political Discourse: Imbalance in State of the Union Addresses 0.20 -0.09 96
Critical Discourse Analysis of Media Portrayal on Trump’s Governance -0.01 -0.10 91
Table 7: Interpretation of Top and Bottom 5 Latent Topics on FlipBias.